SF_{6}: THE FORBIDDEN BAND UNVEILED

\underline{V. BOUDON}, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 6303 CNRS–Université de Bourgogne, 9 Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; L. MANCERON, Laboratoire de Dynamique, Interactions et Réactivité, CNRS UMR 7075, 4 Place Jussieu, F-75252 Paris Cedex, France; F. KWABIA-TCHANA, Laboratoire Interuniversitaire des Systèmes Atmosphériques, UMR CNRS 7583, Université Paris-Est Créteil et Université Paris-Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil Cedex, France; P. ROY, Ligne AILES – Synchrotron SOLEIL, L’Orme des Merisiers, F-91192 Gif-sur-Yvette, France.

Sulfur hexafluoride (SF_{6}) is a greenhouse gas of anthropogenic origin, whose strong infrared absorption in the ν_{3} S–F stretching region near 948 cm^{−1} induces a global warming potential 23900 times bigger than CO_{2}. This heavy species features many hot bands at room temperature (at which the ground state population is only 30 %), especially those originating from the v_{6} = 1 state. Unfortunately, the ν_{6} band itself (near 347 cm^{−1}) being, in first approximation, both infrared and Raman inactive, no reliable information could be obtained about it up to now. A long time ago, some authors suggested^{a} that this band may be slightly activated through Coriolis interaction and may appear as a very faint band, with an integrated intensity about 2 millionths of that of ν_{3}. Using a new cryogenic multipass cell with 93 m optical path length and regulated at 165 ± 2 K temperature, we recorded a spectrum of the ν_{6} far-infrared region thanks to the performances of the AILES Beamline at the SOLEIL french synchrotron facility. Low temperature was used to avoid the presence of the 2ν_{6} − ν_{6} hot band and to reduce the neighboring, stronger ν_{4} − ν_{2} difference band. We are thus able to confirm that the small feature in this region, previously viewed at low-resolution is indeed ν_{6}. We present its fully resolved spectrum. It appears to be activated thanks to unidentified faint interactions resulting in the presence of a first-order dipole moment term that induces unusual selection rules. This spectrum was analyzed thanks to the XTDS software package^{b}, leading to accurate molecular spectroscopic parameters that should be useful to model the hot bands of SF_{6}.

aW. B. Person, B. J. Krohn, J. Mol. Spectrosc. 98, 229–257 (1983), C. Chappados, G. Birnbaum, J. Mol. Spectrosc. 105, 206–214 (1984).

bCh. Wenger, V. Boudon, M. Rotger, M. Sanzharov and J.-P. Champion, J. Mol. Spectrosc., 251 102–113 (2008).